OpenStudy (anonymous):

solve the quadratic equations x2 +3x+0=0

OpenStudy (campbell_st):

all you need to do... if its written correctly is factor \[x^2 + 3x = 0... so ....x(x + 3) = 0\] that should make it easier

OpenStudy (anonymous):

thats the answer wow

OpenStudy (campbell_st):

nope... you need to find the values of x that make the equation try... one answer is x = 0 the other is found by solving x + 3 = 0

OpenStudy (anonymous):

x=0,-3

OpenStudy (anonymous):